# Shear viscosity of quark-gluon plasma at finite temperature and chemical potential and QCD phase transitions

### v1

@@ Abstract
We explore the shear viscosity of quark-gluon plasma (QGP) in the full QCD phase diagram within the framework of kinetic theory with the relaxation time approximation based on the $2\leftrightarrow 2$ elastic scatterings of quark quasiparticles. The temperature and chemical potential dependent masses of particles, including $u,d,s$ quarks, their antiparticles, and exchanged mesons, are calculated in the Polyakov-loop extended Nambu–Jona Lasinio (PNJL) model. The results indicate that, at small chemical potential, the value of $\eta/s$ has a minimum near the Mott dissociation of mesons and increases rapidly in the lower-temperature side of the chiral crossover phase transition. At large chemical potential (high density), $\eta/s$ in the QGP phase is dominated by the temperature, and the value of $\eta/s$ is greatly enhanced at lower temperature. At intermediate temperature and chemical potential near the QCD phase transition, the situation is relatively complicated. The behavior of $\eta/s$ is influenced by the competition between temperature, density effect, and QCD phase transition.
@@
Understanding the nature of quark-gluon plasma and the phase transition to hadronic matter is presently one of the most concerned questions in quantum chromodynamics (QCD). The calculation from lattice QCD and hadron resonance gas model indicates that the QCD phase transition is a smooth crossover at small chemical potential Aoki et al. (2006); Gupta et al. (2011); Bazavov et al. (2012); Borsányi et al. (2013); Bazavov et al. (2014, 2017); Borsányi et al. (2014); Borsanyi et al. (2020). The simulation of lattice QCD is limited at large baryon chemical potential due to the fermion-sign problem. Many QCD inspired approaches/models predict that a first-order phase transition with a second-order critical end point (CEP) exists at finite temperature and chemical potential, connecting with the crossover transformation (e.g., Fukushima (2008); Ratti et al. (2006); Meisinger and Ogilvie (1996); Costa et al. (2010); Sasaki et al. (2012); Ferreira et al. (2014); Schaefer et al. (2010); Skokov et al. (2011); Qin et al. (2011); Gao et al. (2016); Fischer et al. (2014); Maslov and Blaschke (2023); Fu et al. (2020); Shao et al. (2018); He et al. (2022); Liu et al. (2021)).
@@
The theoretical description of the dynamical evolution of QGP in heavy-ion collisions is needed to build the connection with experimental data. The relativistic viscous hydrodynamics provides an effective framework to describe the collective behaviors and transport properties of QGP Israel and Stewart (1979); Baier et al. (2008); Denicol et al. (2012); Jeon and Heinz (2015). The shear $\eta$ and bulk $\zeta$ viscosities are parameters that quantify the dissipation processes in the evolution of fluid. They reflect the deviation of a medium from an ideal fluid. The simulations indicate that the final hadronic observables are sensitive to the two viscous parameters Romatschke and Romatschke (2007); Song and Heinz (2008); Denicol et al. (2009).
@@
In the earlier studies, the ratios of shear and bulk viscosity to entropy density $\eta/s$ and $\zeta/s$ were taken as constants to simulate the evolution of QGP Heinz and Snellings (2013); Gale et al. (2013). One crucial achievement from the combination of heavy-ion collisions and the viscous hydrodynamics was the discovery that the QGP is a nearly perfect fluid Teaney (2003); Shuryak and Zahed (2004); Shuryak (2005) with the shear viscosity to entropy density ratio $\eta/s$ at very small chemical potential close to $1/{4\pi}$, the KSS limit derived in anti–de Sitter space/conformal field theory (AdS/CFT) correspondence Kovtun et al. (2005); Policastro et al. (2001). Subsequently, it was argued that the value of $\eta/s$ is sensitive to the phase transitions of QGP medium. Similar to water, helium and nitrogen, the shear viscosity to entropy ratio $\eta/s$ of QGP exhibits a minimum at the transition from hadrons to quarks and gluons Csernai et al. (2006); Lacey et al. (2007). Recently, the Bayesian statistical analysis has been adopted to constrain the temperature dependent behavior of $\eta/s$ and $\zeta/s$ at vanishing or small chemical potential Bernhard et al. (2019, 2016); Everett et al. (2021); Nijs et al. (2021); Shen et al. (2023).
The beam energy scan (BES) program at RHIC with the collision energy from 3 to 200 GeV provides a great opportunity to explore the evolution and phase transition of QGP in a wide range of temperature and baryon chemical potential Luo and Xu (2017). One important goal of the BES program is to look for the CEP and the boundary of the first-order chiral phase transition at finite temperature and chemical potential. There are also other planned projects to explore the phase boundary of strongly interacting matter at lower temperature and higher density at the Facility for Antiproton and Ion Research (FAIR) in Darmstadt, the Japan Proton Accelerator Research Complex (J-PARC), the Nuclotron-based Ion Collider Facility (NICA) in Dubna and the High Intensity heavy-ion Accelerator Facility (HIAF) in Huizhou.
The hydrodynamic simulations of the evolution of hot and dense QGP in the corresponding energy regime will play an important role in the analysis of hadronic observables in experiments, and will be helpful to diagnose the QCD phase transition. As the major input parameters, the $T$ and $\mu$ dependent $\eta/s$ and $\zeta/s$ are required in the simulation of hydrodynamical evolution of QGP. Since the direct calculation of these transport coefficients from the non-perturbative QCD on lattice is difficult Kovtun et al. (2005); Astrakhantsev et al. (2017), a growing number of studies have explored them in effective models at vanishing or small chemical potential, such as the hadron resonance gas (HRG) model Noronha-Hostler et al. (2009); Tiwari et al. (2012); Noronha-Hostler et al. (2012); Kadam and Mishra (2015), color string percolation model Sahu and Sahoo (2021), linear sigma model Heffernan et al. (2020), quasiparticle models Mitra and Chandra (2017); Moreau et al. (2019); Soloveva et al. (2020); Mykhaylova et al. (2019); Mykhaylova and Sasaki (2021), Green–Kubo formalism Plumari et al. (2012); Kharzeev and Tuchin (2008); Harutyunyan and Sedrakian (2017); Czajka and Jeon (2017); Gubser et al. (2008); Li et al. (2015), transport approach Wesp et al. (2011); Ozvenchuk et al. (2013); Rose et al. (2018), and QCD motivated model Haas et al. (2014); Christiansen et al. (2015); Dubla et al. (2018); Ghiglieri et al. (2018); Danhoni and Moore (2023); Gao and Liu (2018); Zhang and Hou (2018); Chen et al. (2010, 2013); Gao and Liu (2018).
With the progress of BES program at RHIC, more attentions are paid on the transport coefficients of QGP at finite densities (e.g., Reichert et al. (2021); Shen et al. (2023); Deng et al. (2023) and references therein). To extract the transport coefficients at finite chemical potential where a first-order phase transition and/or the critical region may be involved, a feasible method is taking the calculation in effective quark models describing well both the chiral and confinement phase transitions of QCD. In Ref. Sasaki and Redlich (2010), the shear and bulk viscosities of QGP at finite temperature and chemical potential were derived in the two-flavor quark model. Then the shear viscosity and electric conductivity of hot and dense QGP were calculated in the three flavor quark model Marty et al. (2013); Soloveva et al. (2021). In Ref. McLaughlin et al. (2022), to study the shear viscosity in the presence of a crossover phase transition at finite chemical potential, a combination of the interacting hadron resonance gas in the confined phase and the QCD-based assumption in the deconfined phase was constructed McLaughlin et al. (2022); Borsányi et al. (2018). Note that there are also studies in the NJL-type model with a focus on zero or small chemical potential (e.g., Zhuang et al. (1995); Rehberg et al. (1996a); Xiao et al. (2014)). For more studies related to transport coefficients at finite chemical potential, one can also refer to the researches (e.g., Deb et al. (2016); Soloveva et al. (2020); Kadam et al. (2021)).
An important issue that deserves further exploration is the viscosity of QGP at finite temperature and density as well as its relation with the QCD phase transition. In this work, we will elaborate on the shear viscosity of QGP in the full QCD phase diagram in the framework of three-flavor PNJL quark model. The calculation will be performed within the kinetic theory with the relaxation time approximation based on the $2\leftrightarrow 2$ elastic scatterings between $u,d,s$ quarks and their antiparticles. To derive the cross section and flavor dependent relaxation time, the medium dependent masses of quasiparticles will be calculated. The relationship between $\eta/s$ and the meson Mott dissociation as well as the chiral phase transition, including the critical region and the first-order phase transition with the associated spinodal structure, will be explored in detail.
@@
The paper is organized as follows. In Sec. II, we introduce the 2+1 flavor PNJL quark model and the formulas to calculate the shear viscosity in kinetic theory with the relaxation time approximation. In Sec. III, we illustrate the numerical results of cross section, flavor-dependent relaxation time, the shear viscosity to entropy density ratio, and discuss their relations with the QCD phase transitions. A summary is given in Sec. IV.
@@
II three-flavor PNJL quark model and formula of shear viscosity
@@ II.1 The 2+1 flavor PNJL quark model
[…]
@@
where $\psi$ denotes the quark fields with three flavors, $u,\ d$, and $s$; $\hat{m}_{0}=\texttt{diag}(m_{u},\ m_{d},\ m_{s})$ in flavor space; $G$ and $K$ are the four-point and six-point interacting constants, respectively. The $\hat{\mu}=diag(\mu_{u},\mu_{d},\mu_{s})$ are the quark chemical potentials. The covariant derivative is defined as $D_{\mu}=\partial_{\mu}-iA_{\mu}$. The gluon background field $A_{\mu}=\delta_{\mu}^{0}A_{0}$ is supposed to be homogeneous and static, with $A_{0}=g\mathcal{A}_{0}^{\alpha}\frac{\lambda^{\alpha}}{2}$, where $\frac{\lambda^{\alpha}}{2}$ is $SU(3)$ color generators.
@@
[…]
where $\beta=1/T$ is the inverse of temperature and $A_{4}=iA_{0}$. The Polyakov-loop effective potential Roessner et al. (2007) taken in this study is
@@
The mesons are constructed with the quark-antiquark effective interaction. The pseudoscalar mesons $(\pi^{0},\pi^{\pm},K^{0},\overline{K^{0}},K^{\pm},\eta,\eta^{\prime})$, and the scalar meson partners $\left(\sigma_{\pi^{0}},\sigma_{\pi^{\pm}},\sigma_{K^{0}},\sigma_{\overline{K^{0}}},\sigma_{K^{\pm}},\sigma,\sigma^{\prime}\right)$ are considered to describe the elastic scatterings between quarks and antiquarks. In Eq. (1) the six-fermion interaction in the determinant term can be reduced to an effective four-point interaction in the mean-field approximation, by contracting out $\bar{\psi}\psi$ pairs. For $u$ and $d$ quarks, we currently consider the $SU(2)$ isospin symmetry with the degeneracy condition $m_{u}=m_{d}$ and $\mu_{u}=\mu_{d}$, the effective Lagrangian can be then derived as
$\mathcal{L}=$ $\bar{\psi}(i\gamma^{\mu}D_{\mu}\!+\!\gamma_{0}\hat{\mu}\!-\!\hat{m}_{0})\psi-U(\Phi[A],\bar{\Phi}[A],T)$ (6)
@@
$+K_{08}^{+}[\left(\bar{\psi}i\gamma_{5}\lambda^{8}\psi\right)\left(\bar{\psi}i\gamma_{5}\lambda^{0}\psi\right)$
$\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,+\left(\bar{\psi}i\gamma_{5}\lambda^{0}\psi\right)\left(\bar{\psi}i\gamma_{5}\lambda^{8}\psi\right)].$
@@
[…]
With the isospin symmetry of $u$ and $d$ quark, the $\pi^{0}$ is decoupled from the $\eta$ and $\eta^{\prime}$, and the $\pi^{\pm},\pi^{0}$ are degenerate. Similarly, $K_{4}^{\pm}=K_{6}^{\pm}$, which means that the neutral and charged kaons have the same mass. The quark-quark scattering amplitude will be calculated in the random phase approximation, which leads to
@@
$01-4K_{3}^{+}\Pi_{u\bar{u}}^{P}\left(M_{\pi}-i\frac{\Gamma_{\pi}}{2},0\right)=$ 0 (14)
@@
where $\Gamma_{\pi}$ and $\Gamma_{K}$ are the decay width of pion and kaon, respectively. With the increase of temperature, when the value of $\Gamma_{\pi}$ or $\Gamma_{K}$ changes from zero to nonzero, the dissociation from mesons to quarks occurs, which is also called the Mott phase transition. The polarization function $\Pi_{f\bar{f}^{\prime}}^{P(S)}\left(p_{0},\vec{p}\right)$ is
@@
×$\left.\times B_{0}\left(|\mathbf{p}|,M_{f},\mu_{f},M_{f^{\prime}},\mu_{f^{\prime}},p_{0},T\right)\right\},$
@@
In the numerical calculation, a cut-off $\Lambda$ is implemented in the three-momentum space for divergent integrations. We take the model parameters obtained in Rehberg et al. (1996b): $\Lambda=602.3$ MeV, $G\Lambda^{2}=1.835$, $K\Lambda^{5}=12.36$, $m_{u,d}=5.5$ and $m_{s}=140.7$ MeV, determined by fitting $f_{\pi}=92.4$ MeV, $M_{\pi}=135.0$ MeV, $m_{K}=497.7$ MeV and $m_{\eta}=957.8$ MeV.
II.2 Shear viscosity in kinetic theroy with the relaxation time approximation
We shall calculate the transport coefficients in the PNJL model based on the kinetic theory with the relaxation time approximation. The quarks, antiquarks and mesons are taken as quasiparticle in the elastic scatterings. It has been found that the transport coefficients derived in kinetic theory with the quasiparticle approximation are close to the one-loop diagram calculations based on Kubo relations. In this study, we will focus on the properties of shear viscosity at different temperature and baryon density.
The shear viscosity from the kinetic theory with relaxation time approximation Hosoya and Kajantie (1985); Gavin (1985); Sasaki and Redlich (2009, 2010); Chakraborty and Kapusta (2011); Albright and Kapusta (2016) are written as
$\eta=\frac{1}{15T}\sum_{\alpha}\int\frac{d^{3}p}{(2\pi)^{3}}\frac{p^{4}}{E_{\alpha}^{2}}d_{q}\left(\tau_{\alpha}f_{\alpha}^{0}(1-f_{\alpha}^{0})\right)$ (21)
where the subscript $\alpha$ stands for quarks ($u,d,s$) and antiquarks ($\bar{u},\bar{d},\bar{s}$) and the degeneracy factor $d_{q}=2N_{c}$ for spin and color. The relaxation time $\tau_{\alpha}$ for particles $\alpha$ depends on the environment of QGP medium, i.e., it varies with temperature and chemical potential. $f^{0}_{\alpha}\left(E_{\alpha},T,\mu_{\alpha}\right)$ describes the distribution function in local equilibrium for quarks and antiquarks. The effective distribution function given in Eqs. (10) and  (11) will be used in the specific calculation.
Eq. (21) indicates that the relaxation time plays a very important role in calculating the shear viscosity coefficient. It is related to the microscopic scattering process between particles in the thermal medium. In this work, the two-body elastic scatterings, including the quark-quark, quark-antiquark and antiquark-antiquark, are considered to calculate the shear viscosity of QGP. The averaged relaxation time of particle species $i$ can be derived as

### v2

# Shear viscosity and electric conductivity of quark matter at finite temperature and chemical potential with QCD phase transitions

@@ Abstract
In the Beam Energy Scan phase II (BES-II) experiments at RHIC STAR, the quark-gluon plasma (QGP) produced with changing collision energies may probe different regions of the QCD phase diagram. Correspondingly, studying the transport coefficients of quark matter in these regions will contribute to extracting the QCD phase structure through hydrodynamic approaches. We investigate the shear viscosity and electric conductivity within the framework of kinetic theory with the relaxation time approximation, in particular their behaviors near the Mott and first-order phase transitions with a spinodal structure as well as along the isentropic trajectories. To derived the scattering cross-section under different conditions, the temperature and chemical potential dependent masses of quarks, antiquarks and exchanged mesons are calculated in the Polyakov-loop extended Nambu–Jona Lasinio (PNJL) model. The numerical results indicate that, at small chemical potential, the shear viscosity to entropy density ratio ($\eta/s$) has a minimum near the Mott phase transition and increases rapidly in the lower-temperature side of the chiral crossover phase transition. At large chemical potential (high baryon density), $\eta/s$ in the QGP phase is dominated by temperature, and the value of $\eta/s$ is greatly enhanced at low temperatures. At intermediate temperature and chemical potential near the QCD phase transition, the behavior of $\eta/s$ is influenced by the competition between temperature, density, and QCD phase transition. The electirc conductivity ($\sigma/T$) roughly exhibits similar characteristics to $\eta/s$ in the QCD phase diagram, whereas the dimensionless ratio of $\eta/s$ to $\sigma/T$ decreases monotonically with growing temperature, approaching a constant in the high-temperature limit.
@@
Understanding the nature of quark-gluon plasma and the phase transition to hadronic matter is currently one of the most important questions in quantum chromodynamics (QCD). The calculation from lattice QCD and hadron resonance gas model indicates that the QCD phase transition is a smooth crossover at small chemical potential [1, 2, 3, 4, 5, 6, 7, 8]. The simulation of lattice QCD is limited at large baryon chemical potential due to the fermion-sign problem. Many QCD inspired approaches/models predict that a first-order phase transition with a second-order critical end point (CEP) exists at finite temperature and chemical potential, connecting with the crossover transformation (e.g., [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]).
@@
The theoretical description of the dynamical evolution of QGP in heavy-ion collisions is needed to build the connection with experimental data. The relativistic viscous hydrodynamics provides an effective framework to describe the collective behaviors and transport properties of QGP [25, 26, 27, 28]. The shear viscosity ($\eta$) and bulk viscosity ($\zeta$) are parameters that quantify the dissipation processes in the evolution of fluid. They reflect the deviation of a medium from an ideal fluid. The simulations indicate that the final hadronic observables are sensitive to the two viscous parameters [29, 30, 31].
@@
In the earlier studies, the ratios of shear and bulk viscosity to entropy density $\eta/s$ and $\zeta/s$ were taken as constants to simulate the evolution of QGP in high-energy collisions [32, 33]. One crucial achievement from the combination of heavy-ion collisions and the viscous hydrodynamics was the discovery that the QGP is a nearly perfect fluid [34, 35, 36] with the shear viscosity to entropy density ratio $\eta/s$ at very small chemical potential close to $1/{4\pi}$, the KSS limit derived in anti–de Sitter space/conformal field theory (AdS/CFT) correspondence [37, 38]. Subsequently, it was argued that the value of $\eta/s$ is sensitive to the phase transitions of QGP medium. Similar to water, helium and nitrogen, the shear viscosity to density ratio $\eta/s$ of QGP possibly exhibits a minimum at the transition from hadrons to quarks and gluons [39, 40]. Since the direct calculation of these transport coefficients from the non-perturbative QCD on lattice is difficult [37, 41], a growing number of studies were performed in effective models, such as the hadron resonance gas (HRG) model [42, 43, 44, 45], color string percolation model [46], linear sigma model [47], quasiparticle models [48, 49, 50, 51, 52], Green–Kubo formalism [53, 54, 55, 56, 57, 58], transport approach [59, 60, 61], and QCD motivated models [62, 63, 64, 65, 66, 67, 68, 69, 70, 67, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81]. The Bayesian statistical analysis has also been adopted to constrain the temperature dependent behavior of $\eta/s$ and $\zeta/s$ at vanishing or small chemical potential [82, 83, 84, 85, 86].
In experiments, the BES II program at RHIC STAR provides a great opportunity to explore the evolution of QGP in a wide range of temperature and baryon chemical potential [87, 88, 89]. Especially, the critical region and the first-order phase transition may be involved in the BES II experiments with the collision energy declined to several GeV. As the shear viscosity is closely correlated with the collective flows of final state particles, its value are crucial to diagnose the QCD phase transition in future hydrodynamical simulations of QGP evolution with a first-order phase transition. Although some calculations have been performed to explore the shear viscosity of hot and dense quark matter [72, 73, 74, 50, 86, 77, 78, 80], even including the chiral phase transition in the temperature-baryon chemical ($T-\mu_{B}$) plane [72, 73, 78], an issue that needs further investigation is the behavior of shear viscosity of quark matter near the Mott phase transition and the first-order chiral phase transition associated with a spinodal structure as well as the evolution of $\eta/s$ on the isentropic trajectories. On the other hand, the electric conductivity ($\sigma$) is the parameter to quantify the Ohmic response of QGP to an electric field. It is of fundamental importance for the evolution and lifetime of the electromagnetic fields in the QGP medium [48]. In this work, we will elaborate on the shear viscosity and electric conductivity of quark matter in the full QCD phase diagram. The behaviors of $\eta/s$ and $\sigma/T$ on different phase planes ($T-\mu_{B}$ and $T-\rho_{B}$ planes, as well as along the isentropic trajectories) will be explored in detail, particularly their behaviors near the meson Mott dissociation and the chiral phase transition, including the crossover, critical and first-order phase transition with the spinodal structure connected with the metstable and unstable phase. We will also calculate the ratio of $\eta/s$ to $\sigma/T$, and analyze its characteristics in high-temperature limit and near the QCD phase transitions.
@@
The calculations will be performed within the kinetic theory with the relaxation time approximation based on the $2\rightarrow 2$ elastic scatterings between $u,d,s$ quarks and their antiparticles. To derive the cross section and flavor dependent relaxation time, the medium dependent masses of quasiparticles will be calculated in the PNJL model. The paper is organized as follows. In Sec. II, we introduce the 2+1 flavor PNJL quark model and the formulas to calculate the shear viscosity and electric conductivity in kinetic theory with the relaxation time approximation. In Sec. III, we illustrate the numerical results of cross section, flavor-dependent relaxation time, shear viscosity and electric conductivity under different physical coditions, and discuss their relations with the QCD phase transitions. A summary is given in Sec. IV.
@@
II PNJL quark model and formulae of shear viscosity and electric conductivity
@@ II.1 The 2+1 flavor PNJL quark model
$\!\mathcal{L}$ $\!=$ $\!\bar{\psi}(i\gamma^{\mu}\!D_{\mu}\!+\!\gamma_{0}\hat{\mu}\!-\!\hat{m}_{0})\psi\!+\!G\!\sum_{k=0}^{8}\!\big[(\bar{\psi}\lambda_{k}\psi)^{2}\!+\!(\bar{\psi}i\gamma_{5}\lambda_{k}\psi)^{2}\big]$ (1)
@@
where $\psi$ denotes the quark fields with three flavors, u,du,\ d, and $s$; $\hat{m}_{0}=\texttt{diag}(m_{u},\ m_{d},\ m_{s})$ in flavor space; $G$ and $K$ are the four-point and six-point interacting constants, respectively. The $\hat{\mu}=diag(\mu_{u},\mu_{d},\mu_{s})$ are the quark chemical potentials. The covariant derivative is defined as $D_{\mu}=\partial_{\mu}-iA_{\mu}$. The gluon background field $A_{\mu}=\delta_{\mu}^{0}A_{0}$ is supposed to be homogeneous and static, with $A_{0}=g\mathcal{A}_{0}^{\alpha}\frac{\lambda^{\alpha}}{2}$, where $\frac{\lambda^{\alpha}}{2}$ is $SU(3)$ color generators.
@@
$L(\vec{x})=\mathcal{P}exp\bigg[i\int_{0}^{\beta}d\tau A_{4}(\vec{x},\tau)\bigg],$ (2)
where $\beta=1/T$ is the inverse of temperature and $A_{4}=iA_{0}$. The Polyakov-loop effective potential [90] taken in this study is
@@
The mesons are constructed with the quark-antiquark effective interaction. The pseudoscalar mesons $(\pi^{0},\pi^{\pm},K^{0},\overline{K^{0}},K^{\pm},\eta,\eta^{\prime})$, and the scalar meson partners $\left(\sigma_{\pi^{0}},\sigma_{\pi^{\pm}},\sigma_{K^{0}},\sigma_{\overline{K^{0}}},\sigma_{K^{\pm}},\sigma,\sigma^{\prime}\right)$ are considered to describe the elastic scatterings between quarks and antiquarks. In Eq. (1) the six-fermion interaction in the determinant term can be reduced to an effective four-point interaction in the mean-field approximation, by contracting out $\bar{\psi}\psi$ pairs. We currently consider the $SU(2)$ isospin symmetry for $u$ and $d$ quarks with the degeneracy condition $m_{u}=m_{d}$ and $\mu_{u}=\mu_{d}$, the effective Lagrangian can be then derived as
$\mathcal{L}=$ $\bar{\psi}(i\gamma^{\mu}D_{\mu}\!+\!\gamma_{0}\hat{\mu}\!-\!\hat{m}_{0})\psi-U(\Phi[A],\bar{\Phi}[A],T)$
@@
$+K_{08}^{+}[\left(\bar{\psi}i\gamma_{5}\lambda^{8}\psi\right)\left(\bar{\psi}i\gamma_{5}\lambda^{0}\psi\right)+\left(\bar{\psi}i\gamma_{5}\lambda^{0}\psi\right)\left(\bar{\psi}i\gamma_{5}\lambda^{8}\psi\right)].$
@@
where $E_{i}=\sqrt{\mathbf{p}^{2}+M_{i}^{2}}$ is the dispersion relation of (anti)quarks in QCD medium.
With the isospin symmetry of $u$ and $d$ quark, the $\pi^{0}$ is decoupled from the $\eta$ and $\eta^{\prime}$, and the $\pi^{\pm},\pi^{0}$ are degenerate. Similarly, $K_{4}^{\pm}=K_{6}^{\pm}$, which means that the neutral and charged kaons have the same mass. The quark-quark scattering amplitude will be calculated in the random phase approximation [80], which leads to
@@
$1-4K_{3}^{+}\Pi_{u\bar{u}}^{P}\left(M_{\pi}-i\frac{\Gamma_{\pi}}{2},0\right)=0$ (14)
@@
where $\Gamma_{\pi}$ and $\Gamma_{K}$ are the decay width of pion and kaon, respectively. With the increase of temperature, when the value of $\Gamma_{\pi}$ or $\Gamma_{K}$ changes from zero to nonzero, the dissociation from mesons to quarks occurs, which is also called the Mott phase transition. The polarization function $\Pi_{f\bar{f}^{\prime}}^{P(S)}\left(p_{0},\vec{p}\right)$ derived in Ref. [80] is
@@
$\left.\times B_{0}\left(|\mathbf{p}|,M_{f},\mu_{f},M_{f^{\prime}},\mu_{f^{\prime}},p_{0},T\right)\right\},$
@@
In the numerical calculation, a cut-off $\Lambda$ is implemented in the three-momentum space for divergent integrations. We take the model parameters obtained in [91]: $\Lambda=602.3$ MeV, $G\Lambda^{2}=1.835$, $K\Lambda^{5}=12.36$, $m_{u,d}=5.5$ and $m_{s}=140.7$ MeV, determined by fitting $f_{\pi}=92.4$ MeV, $M_{\pi}=135.0$ MeV, $m_{K}=497.7$ MeV and $m_{\eta}=957.8$ MeV.
II.2 Shear viscosity and electric conductivity in kinetic theroy with the relaxation time approximation
We shall calculate the transport coefficients in the PNJL model based on the kinetic theory in which quarks, antiquarks and mesons are taken as quasiparticle. The shear viscosity and electric conductivity from the kinetic theory with relaxation time approximation[92, 93, 94, 71, 95, 96, 50] are written as
$\eta=\frac{1}{15T}\sum_{\alpha}d_{q}\tau_{\alpha}\int\frac{d^{3}p}{(2\pi)^{3}}\frac{p^{4}}{E_{\alpha}^{2}}f_{\alpha}^{0}(1-f_{\alpha}^{0})$ (21)
$\sigma_{e}=\frac{e^{2}}{3T}\sum_{\alpha}q_{\alpha}^{2}d_{q}\tau_{\alpha}\int\frac{d^{3}p}{(2\pi)^{3}}\frac{p^{2}}{E_{\alpha}^{2}}f_{\alpha}^{0}(1-f_{\alpha}^{0})$ (22)
where the subscript $\alpha$ stands for quarks ($u,d,s$) and antiquarks ($\bar{u},\bar{d},\bar{s}$), $d_{q}=2N_{c}$ is the degeneracy factor for spin and color, $e^{2}=4\pi/137$ is the scaled charge-squared with quark charges $q_{i}=+2/3(u),-1/3(d),-1/3(s)$. The relaxation time $\tau_{\alpha}$ for particles $\alpha$ depends on the environment of QGP medium, i.e., it varies with temperature and chemical potential. $f^{0}_{\alpha}\left(E_{\alpha},T,\mu_{\alpha}\right)$ describes the distribution function in local equilibrium for quarks and antiquarks. The effective distribution function given in Eqs. (10) and  (11) will be used in the calculation.
Eq. (21) indicates that the relaxation time plays a very important role in calculating the shear viscosity coefficient. It is related to the microscopic scattering process between particles in the thermal medium. In this work, the $2\rightarrow 2$ elastic scatterings, including the quark-quark, quark-antiquark and antiquark-antiquark, are considered to calculate the shear viscosity of QGP. The averaged relaxation time of particle species $i$ can be derived as